OpenStudy (anonymous):

What is the equation of the line that goes through (2,4) and (3,-1) ?

OpenStudy (anonymous):

first find the slope, do you know how to get the slope?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do you do after?

OpenStudy (anonymous):

okay, well once you have it, we will use m to represent slope.

OpenStudy (anonymous):

then plug it into this: y-y1=m(x-x1)

OpenStudy (anonymous):

Plug that slope along with one of your points into the point slope formula: \[y - y_1 = m(x-x_1)\] Where m is the slope you found, and (x1, y1) is either of your two points.

OpenStudy (anonymous):

another formula (x-2)/(y-4)=(2-3)/(4+1)